# Improving accuracy of tree-tensor network approach by optimization of network structure

### v1

@@ III.1 Types of calculation and Quantities computed
We have applied the algorithm discussed in the previous section to the random XY-exchange model under random magnetic field and the Richardson model. In the calculation, we can choose (i) the initial TTN, (ii) the scheme to update the local TTN structure, and (iii) the value of $\chi$ with which the structural optimization is performed. We have carried out various types of calculations specified by (i), (ii), and (iii) and compared them to evaluate their performance. The types of calculations performed are in the following.
@@
    • (1)
      MPN: We prepared the MPN as the initial TTN. The bare spins were arranged one-dimensionally in the order of the site index and the canonical center is located at the center of the MPN, as shown in Fig. 1 (a).
    • (2)
      PBT: We prepared the PBT as the initial TTN. The bare spins were arranged one-dimensionally in the order of the site index and the canonical center is located at the “top” of the PBT, as shown in Fig. 1 (b).
    • (3)
      TTN constructed by tSDRG method: We prepared the initial TTN using the tSDRG, which is a method to construct a TTN suitable to represent the ground state of a system with randomness. See Ref. [29] for the details. We adopted the gap $\Delta^{\rm I}_{\rm max}$ defined in Ref. [29] to determine the order of the isometries to be merged. This initial TTN constructed by tSDRG was used only for the random XY model.
@@
    • (A)
      No structural optimization: The TTN structure was fixed to the one of the initial TTN, and only the isometries were improved. This calculation is equivalent to the conventional variational TTN calculation.
    • (B)
@@
    • (D)
      Stochastic selection using the probability $P^{(1)}$: We optimized the TTN structure by selecting the local connection stochastically according to the probability defined by Eq. (5). For the effective inverse temperature, we employed the exponential form,
      $\beta=\beta_{0}~{}2^{n/n_{\tau}}$ (7)
@@
    • (E)
      Stochastic selection using the probability $P^{(2)}$: We optimized the TTN structure by selecting the local connection stochastically according to the probability defined by Eq. (6). The effective inverse temperature was treated in the same way as in (D).
@@
    In schemes (B)-(E) in (ii), we first performed the structural optimization calculation with $\chi=\chi_{\rm opt}$ to determine the optimal TTN structure. After that, the isometries were further improved with increasing $\chi$ in the TTN with the structure fixed to the optimized one. In the actual calculation for the random XY model (the Richardson model), we performed the structural optimization at $\chi_{\rm opt}=16$ or $32$ ($20$ or $40$), and then continued the calculation to optimize isometries with increasing $\chi$ up to $40$ ($80$). The details, including the number of sweeps and the criteria for judging convergence are given in Appendix B.
At each step of the calculation, we computed the lowest energy eigenvalue $E_{0}$ of the effective Hamiltonian $\tilde{\mathcal{H}}$ constructed for the central area and the EE on the center bond. We then adopted the lowest $E_{0}$ obtained in a sweep as the variational energy at the sweep. Here, it should be noted that $E_{0}$ is the eigenvalue of the effective Hamiltonian expressed in the Hilbert space expanded by the four bonds $a,b,c$, and $d$ in Fig. 3(b), and the corresponding eigenstate $\Psi(abcd)$, that is expressed as Eq. (2) is not strictly the same as the TTN state in the sense that the center bond does not suffer from the cutoff. Therefore, as a wave function corresponding to the TTN state in which the dimensions of all bonds are upper bounded by $\chi$, we constructed the truncated wave function as
@@
for the case that the local connection $(ab|cd)$ [Fig. 3 (c)] was selected, and by similar ways for the other connections. We then calculate the expectation value of the effective Hamiltonian on the truncated state as
@@
III.2 Random XY-exchange Model Under Random Magnetic Field
We have applied the TTN structural optimization algorithm to the spin-1/2 model with random XY-exchange interactions under random magnetic field. The model Hamiltonian is given by
@@ III.2 Random XY-exchange Model Under Random Magnetic Field
where $J_{\rm r}$ is a parameter to control the strength of the exchange interactions and ${\bm{S}}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the spin-1/2 operator at $i$ th site. $\epsilon_{ij}$ and $h_{i}$ are random variables obeying uniform distributions in the range $[0,1]$ and $\left[-\frac{1}{2},\frac{1}{2}\right]$, respectively. We have performed the calculations for the case of the exchange constant $J_{\rm r}=8$ and the system size $N=64$. In the calculations, we have prepared $\mathcal{N}_{\rm s}=200$ random samples and computed the lowest-energy state in the subspace with zero total magnetization, $\sum_{i}S^{z}_{i}=0$. Note that this state may not be the ground state of each random sample, depending on the random model parameters $\{\epsilon_{ij}\}$ and $\{h_{i}\}$.
@@
We have employed MPN, PBT, and the TTN obtained by tSDRG [(1), (2), and (3) introduced in Sec. III.1] as the initial TTN structure and performed the calculations of schemes (A)-(E). In scheme (A) (without TTN structural optimization), we optimized isometries in the TTN with fixed structure with sequentially increasing $\chi$ from $16$ to $40$. For schemes (B)-(E), two cases of calculations were performed: in one case, the TTN structural-optimization calculation was executed with $\chi=\chi_{\rm opt}=16$, and in the other case, the calculation was done with $\chi=\chi_{\rm opt}=32$. After that, we fixed the TTN structure and optimized only isometries with increasing $\chi$ up to $40$. See Appendix B for details. Thus, the type of calculation is specified by three conditions: initial TTN [(1)-(3)], optimization scheme [(A)-(E)], and $\chi_{\rm opt}$ ($=16$ or $32$) with which the TTN structural optimization was performed. For the stochastic structural optimization [schemes (D) and (E)], we performed two runs of the calculation with the different annealing parameters, $(\beta_{0},n_{\tau})=(0.1,2),(0.1,4)$, in Eq. (7). Then, among the results of the two runs, we adopted the one obtained by the run that gives the smaller variational energy $E_{0}$ ($E^{\prime}_{0}$) at $\chi=40$ as the lowest energy used in the analysis of random average of $\delta r$ ($\delta r^{\prime}$) defined below.
@@
To evaluate the performance of each type of calculation, we explore the relative reduction in the variational energy with respect to the energy obtained by the calculation of the fixed structure of MPN [type (1A)] at $\chi=16$,
$\delta r(\nu,{\rm X},\chi)$ $=$ $\frac{E_{0}(\nu,{\rm X},\chi)-E_{0}(\nu,{\rm 1A},16)}{|E_{0}(\nu,{\rm 1A},16)|},$ (11)
$\delta r^{\prime}(\nu,{\rm X},\chi)$ $=$ $\frac{E^{\prime}_{0}(\nu,{\rm X},\chi)-E^{\prime}_{0}(\nu,{\rm 1A},16)}{|E^{\prime}_{0}(\nu,{\rm 1A},16)|},$ (12)
@@
where $E^{(^{\prime})}_{0}(\nu,{\rm X},\chi)$ is the variational energy obtained by the calculation of type ${\rm X}$ at the bond dimension $\chi$ for the random sample $\nu=1,...,\mathcal{N}_{\rm s}$. We then examine the random average of $\delta r(\nu,{\rm X},\chi)$ and $\delta r^{\prime}(\nu,{\rm X},\chi)$, respectively denoted as $[\delta r({\rm X},\chi)]$ and $[\delta r^{\prime}({\rm X},\chi)]$. ([]delimited-[][$\cdots$][ ⋯ ] denotes the average over $\mathcal{N}_{\rm s}$ random samples.) In the following, we present only the results of $[\delta r({\rm X},\chi)]$ since we have found that $[\delta r({\rm X},\chi)]$ and $[\delta r^{\prime}({\rm X},\chi)]$ lead to essentially the same conclusions.
@@
Figure 4 shows the random average $[\delta r({\rm X},\chi)]$ as a function of $\chi$. Also, the values of $[\delta r({\rm X},\chi=40)]$ for all types are plotted in Fig. 5. The magnitude of the error bar for the random average $[\delta r({\rm X},\chi)]$ shown in the figure is obtained as
$\sigma_{r}=\sqrt{\frac{1}{\mathcal{N}_{\rm s}(\mathcal{N}_{\rm s}-1)}\sum_{\nu=1}^{\mathcal{N}_{\rm s}}\left\{\delta r(\nu,{\rm X},\chi)-[\delta r({\rm X},\chi)]\right\}^{2}}.$
Figure 4: Random averages of the relative reduction in the variational energy, $[\delta r({\rm X},\chi)]$, as functions of $\chi$, for the types where the initial TTN is (a) (1) MPN, (b) (2) PBT, and (c) (3) TTN obtained by tSDRG. In (a), the data for type (1A) are shown in the inset.
@@
Figure 5: Random averages of the relative reduction in the variational energy, $[\delta r({\rm X},\chi)]$, at $\chi=40$. Circles, squares, and triangles represent the results of the types where the initial TTN is (1) MPN, (2) PBT, and (3) TTN obtained by tSDRG, respectively. Solid symbols represent the results of (A) or (B)-(E) with $\chi_{\rm opt}=16$, while open symbols are those of (B)-(E) with $\chi_{\rm opt}=32$.
@@
In the calculation without structural optimization [scheme (A)], (1) MPN has the poorest accuracy in variational energies, whereas (2) PBT and (3) TTN obtained by tSDRG yield the results with relatively good accuracy. As for the results of non-stochastic structural optimization [schemes (B) and (C)], a significant improvement is observed when the initial TTN is MPN, reflecting the poor accuracy of the MPN calculation without structural optimization. When the initial TTN is PBT or that of tSDRG, there is an improvement in accuracy although it is small. Besides, when the initial TTN is MPN and PBT [(1B), (1C), (2B), and (2C)], the accuracy in the variational energies is worse than that of the tSDRG without structural optimization (3A). These results suggest that the structural optimization calculations with schemes (B) and (C) retain the influence of the initial TTN and are trapped in the vicinity of the initial TTN in the TTN-structure space. (Incidentally, this observation indicates the effectiveness of tSDRG for the random XY model.) We note that there is no notable difference between the results of schemes (B) and (C); namely, the performance of the schemes to minimize EE and that to minimize the truncated rDM weight is comparable, at least for the random XY model and $\chi_{\rm opt}$ treated. In the calculations of the stochastic structural optimizations [schemes (D) and (E)], the variational energy is further reduced, yielding the best accuracy for the calculations performed here. Of particular importance is the fact that the results of these schemes achieve almost the same level of accuracy independent of the initial TTN. These results suggest that the stochastic structural optimizations work well for the random XY model and succeed in escaping the influence of the initial TTN. Comparing the results of schemes (D) and (E), scheme (D) yields slightly lower variational energy, if any. However, the difference is small and remains within the error of the random sampling average.
@@
Finally, concerning the effect of $\chi_{\rm opt}$ at which the structural optimization is performed, the results of $\chi_{\rm opt}=32$ yield slightly lower variational energy than those of $\chi_{\rm opt}=16$ in the stochastic structural optimization calculations [schemes (D) and (E)]. However, the difference is within the statistical error and basically smaller than the difference between stochastic optimizations [schemes (D) and (E)] and non-stochastic ones [schemes (B) and (C)]. This observation indicates that the value of $\chi_{\rm opt}$ has only a minor effect comparing the difference in the initial TTN and the updating scheme, at least within the range of $\chi$ we treated here.
@@
Next, we discuss EE on each bond of TTN. Since EE on the physical bonds, the bonds directly connected to the bare spins, takes the same value regardless of the TTN structure, we consider only the EEs on the auxiliary bonds in the following analysis. In the analysis, we use the data of EEs obtained in the sweep which yields the lowest variational energy $E_{0}$ among the sweeps iterated at each $\chi$. For the stochastic schemes [schemes (D) and (E)], we use the EEs obtained by the run that yields the lower variational energy $E_{0}$ at $\chi=40$ between the two runs with different annealing parameters.
@@
Figure 6: (a) $\left[\mathcal{S}_{\rm av}\right]$ and (b) $\left[\mathcal{S}_{\rm max}\right]$, that are respectively the random averages of the average values $\mathcal{S}_{\rm av}(\nu)$ and maximum values $\mathcal{S}_{\rm max}(\nu)$ of EEs on auxiliary bonds in each random sample. Circles, squares, and triangles represent the results at $\chi=40$ of the types where the initial TTN is (1) MPN, (2) PBT, and (3) TTN obtained by tSDRG, respectively. Solid symbols represent the results of (A) or (B)-(E) with $\chi_{\rm opt}=16$, while open symbols are those of (B)-(E) with $\chi_{\rm opt}=32$. Grey crosses represent the results at $\chi=32$ ($\times$) and $\chi=24$ ($+$); the latter presents only for (A) and (B)-(E) with $\chi_{\rm opt}=16$. The results for (1A) are plotted in a different scale for the clarity. Error bars are plotted only for the results at $\chi=40$, although they are smaller than the symbols in most cases.
@@
To see how the EEs in the TTN are suppressed by the structural optimization, we explore the average and maximum values of EE. Figure 6 shows the random averages of $\mathcal{S}_{\rm av}(\nu)$ and $\mathcal{S}_{\rm max}(\nu)$, $\left[\mathcal{S}_{\rm av}\right]$ and $\left[\mathcal{S}_{\rm max}\right]$, where $\mathcal{S}_{\rm av}(\nu)$ and $\mathcal{S}_{\rm max}(\nu)$ are respectively the average and maximum values of EE on the auxiliary bonds in each random sample. Error bars are obtained by a formula similar to Eq. (13). We note that $\left[\mathcal{S}_{\rm av}\right]$ and $\left[\mathcal{S}_{\rm max}\right]$ are increasing functions of $\chi$, and one should analyze them at a sufficiently large $\chi$ to compare the TTNs obtained by different types of structural optimization.
@@
For the average EE, $\left[\mathcal{S}_{\rm av}\right]$, the result for the MPN without structural optimization (1A) is remarkably large and has a pronounced $\chi$ dependence compared to other data, where the convergence with respect to $\chi$ is almost achieved at $\chi=40$. This observation indicates that in the MPN, there is a significant fraction of auxiliary bonds whose EEs are too large to be borne even by a bond with $\chi=40$. The values of $\left[\mathcal{S}_{\rm av}\right]$ of PBT and tSDRG without structural optimization [(2A) and (3A)] are much smaller than that of MPN (1A). The non-stochastic structural optimization [schemes (B) and (C)] reduces $\left[\mathcal{S}_{\rm av}\right]$ for MPN and tSDRG, while no notable reduction is observed for PBT. Furthermore, for all initial TTNs, the values of $\left[\mathcal{S}_{\rm av}\right]$ obtained by schemes (B) and (C) are larger than those of stochastic structural optimization [schemes (D) and (E)]. These observations suggest that the non-stochastic structural optimization by schemes (B) and (C), especially when applied to PBT, is trapped around the initial TTN in the TTN structural space. Finally, the results of $\left[\mathcal{S}_{\rm av}\right]$ obtained by the stochastic structural optimization [schemes (D) and (E)] achieve the smallest values independently of the initial TTNs, although there is some exception of the types (1D) and (1E) with $\chi_{\rm opt}=16$ which yield slightly large values. This result demonstrates the effectiveness of the stochastic optimization for the random XY model. As for the dependence of the results on $\chi_{\rm opt}$, the difference between $\left[\mathcal{S}_{\rm av}\right]$ obtained by each type with $\chi_{\rm opt}=16$ and $\chi_{\rm opt}=32$ is basically small compared to the difference due to the initial TTN or the updating scheme [except for (1D) and (1E)], indicating that the effect of $\chi_{\rm opt}$ is secondary.
@@
The data of $\left[\mathcal{S}_{\rm max}\right]$ suffer from relatively large $\chi$ dependence, however, they lead us to the following conclusions that are basically consistent with the ones we read from the data of $[\mathcal{S}_{\rm av}]$: That is, the results of scheme (A) exhibit specific behaviors reflecting the characteristics of each TTN. [We note that $[\mathcal{S}_{\rm max}]$ for (1A) at $\chi=40$ is smaller than those for other types, but its large $\chi$ dependence implies that $[\mathcal{S}_{\rm max}]$ for (1A) may be greater than those for other types at $\chi\to\infty$.] The results of the non-stochastic schemes [schemes (B) and (C)] exhibit a sizable dependence on the initial TTN. The results of stochastic structural optimization [schemes (D) and (E)] yield the smallest $\left[\mathcal{S}_{\rm max}\right]$ [except (1A)], with small dependence on the initial TTN. In addition, compared to the differences by the initial TTN and structural optimization scheme, the differences by $\chi_{\rm opt}$ are minor.
We also analyze the histogram of EEs on the auxiliary bonds. For each type of calculation, a histogram $P(\mathcal{S})$ is obtained using the values of EEs on the auxiliary bonds in all random samples as the population [the total number of EEs is $\mathcal{N}_{\rm s}(N-3)$], and the bin size is set to $\Delta\mathcal{S}=0.1$. Figure 7 presents the obtained histograms. The $\chi$ dependence is sizable only for (1A) and irrelevant for the other types, which is consistent with the results of $\left[\mathcal{S}_{\rm av}\right]$. The histograms for the scheme without structural optimization [scheme (A)] have a different shape depending on the initial TTNs, which should reflect the characteristics of each TTN. The non-stochastic optimizations applied to PBT [(2B) and (2C)] yield histograms with a similar shape to that of (2A), where the TTN is fixed to be PBT, suggesting that the TTN structures obtained by those types remain within the vicinity of PBT in the TTN structure space and retain the characteristics of PBT, the initial TTN. Also, when the initial TTN is MPN or that of tSDRG, the histograms obtained by schemes (B) and (C) tend to approach those obtained by stochastic optimization [schemes (D) and (E)]. However, they still exhibit notable differences such as bumpy behavior in the region of large EE ($1.2\lesssim\mathcal{S}\lesssim 3.0$). In contrast, the histograms obtained by stochastic optimization [schemes (D) and (E)] converge to nearly the same shape regardless of the initial TTN; $P(\mathcal{S})$ exhibits a sharp peak around $\mathcal{S}\sim 1.0-1.1$, and from the peak, it decreases with a gentle tail as $\mathcal{S}$ increases. These histograms should be the ones of the best TTNs achieved in the present calculations.

### v2

@@ III.1 Types of Calculation and Quantities Computed
We have applied the algorithm discussed in the previous section to the random XY-exchange model under random magnetic fields and the Richardson model. In the calculation, we can choose (i) the initial TTN, (ii) the scheme to update the local TTN structure, and (iii) the value of $\chi$ with which the structural optimization is performed. We have carried out various types of calculations specified by (i), (ii), and (iii) and compared them to evaluate their performance. The types of calculations performed are as follows.
@@
    • $\bullet$
      Matrix-Product Network (MPN): We prepared the MPN as the initial TTN. The bare spins were arranged one-dimensionally in the order of the site index, and the canonical center is located at the center of the MPN, as shown in Fig. 1 (a).
    • $\bullet$
      Perfect-binary tree (PBT): We prepared the PBT as the initial TTN. The bare spins were arranged one-dimensionally in the order of the site index, and the canonical center is located at the “top” of the PBT, as shown in Fig. 1 (b).
    • $\bullet$
      TTN constructed by tSDRG method (tSDRG-TTN): We prepared the initial TTN using the tSDRG, which is a method to construct a TTN suitable to represent the ground state of a system with randomness. See Ref. [29] for the details. We adopted the gap $\Delta^{\rm I}_{\rm max}$ defined in Ref. [29] to determine the order of the isometries to be merged. This initial TTN constructed by tSDRG was used only for the random XY model.
@@
    • $\bullet$
@@
    • $\bullet$
      Stochastic selection: We optimized the TTN structure by selecting the local connection stochastically according to the probability $P^{(1)}$ defined by Eq. (5). For the effective inverse temperature, we employed the exponential form,
      $\beta=\beta_{0}~2^{n/n_{\tau}},$ (7)
@@
    In addition, we have performed the calculation without structural optimization for comparison. In this case, the TTN structure was fixed to that of the initial TTN, and only the isometries were improved. This calculation is equivalent to the conventional variational TTN calculation.
    We note that, for the calculation of the stochastic selection, one may employ the same scheme as the above but using the probability $P^{(2)}$ in Eq. (6) instead of $P^{(1)}$. We have also performed the calculation with $P^{(2)}$ and found that the results of both the stochastic calculations with $P^{(1)}$ and $P^{(2)}$ coincide semiquantitatively. Therefore, in the following arguments, we discuss only the results of the calculation with $P^{(1)}$. The results of the stochastic calculation using $P^{(2)}$ are presented in Appendix C.
@@
    For the calculations with structural optimization, we first performed the structural optimization with $\chi=\chi_{\rm opt}$ to determine the optimal TTN structure. After that, the isometries were further improved by increasing $\chi$ for the TTN with the structure fixed to the optimized one. In the actual calculation for the random XY model (the Richardson model), we performed the structural optimization at $\chi_{\rm opt}=16$ or $32$ (2020 or $40$), and then continued the calculation to optimize isometries by increasing $\chi$ up to $40$ ($80$). The details, including the number of sweeps and the criteria for judging convergence, are given in Appendix B.
At each step of the calculation, we computed the lowest energy eigenvalue $E_{0}$ of the effective Hamiltonian $\tilde{\mathcal{H}}$ constructed for the central area. We then adopted the lowest $E_{0}$ obtained within a sweep as the variational energy at the sweep. Here, it should be noted that $E_{0}$ is the eigenvalue of the effective Hamiltonian expressed in the Hilbert space expanded by the four bonds $a,b,c$, and $d$ in Fig. 3(b), and the corresponding eigenstate $\Psi(abcd)$, that is expressed as Eq. (2) is not strictly the same as the TTN state since the center bond does not suffer from the truncation effect. Therefore, as a wave function corresponding to the TTN state in which the dimensions of all bonds are upper bounded by $\chi$, we constructed the truncated wave function as
@@
for the case that the local connection $(ab|cd)$ [Fig. 3 (c)] was selected, and in similar ways for the other connections. We then calculated the expectation value of the effective Hamiltonian on the truncated state as
@@
III.2 Random XY-exchange Model Under Random Magnetic Fields
We have applied the TTN structural optimization algorithm to the spin-1/2 model with random XY-exchange interactions under random magnetic fields. The model Hamiltonian is given by
@@ III.2 Random XY-exchange Model Under Random Magnetic Fields
where $J_{\rm r}$ is a parameter to control the strength of the exchange interactions and ${\bm{S}}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the spin-1/2 operator at the $i$ th site. $\epsilon_{ij}$ and $h_{i}$ are random variables obeying uniform distributions in the range $[0,1]$ and $\left[-\frac{1}{2},\frac{1}{2}\right]$, respectively. We have performed the calculations for the case of the exchange constant $J_{\rm r}=8$ and the system size $N=64$. In the calculations, we have prepared $\mathcal{N}_{\rm s}=200$ random samples and computed the lowest-energy state in the subspace with zero total magnetization, $\sum_{i}S^{z}_{i}=0$. Note that this state may not be the ground state of each random sample, depending on the random model parameters $\{\epsilon_{ij}\}$ and $\{h_{i}\}$.
@@
We have performed the calculation employing MPN, PBT, and tSDRG-TTN as the initial TTN structures. In the calculation without structural optimization, we have optimized isometries in the TTN with the fixed network structure by sequentially increasing $\chi$ from $16$ to $40$. For the calculations with structural optimization, two cases of $\chi_{\rm opt}$ were tested: we first performed the TTN structural optimization with $\chi=\chi_{\rm opt}=16$ or $32$. Then, we fixed the TTN structure and optimized only isometries by increasing $\chi$ up to $40$. See Appendix B for details. Thus, the type of calculation is specified by three conditions: the initial TTN, optimization scheme, and $\chi_{\rm opt}$ (=16=16 or $32$) with which the TTN structural optimization was performed. For the stochastic structural optimization, we have performed two runs of the calculation with the different annealing parameters, $(\beta_{0},n_{\tau})=(0.1,2),(0.1,4)$, in Eq. (7). Then, among them, we have adopted the result of the run that gives the smaller variational energy $E_{0}$ ($E^{\prime}_{0}$) at $\chi=40$ for the analysis of the random average of $\delta r$ ($\delta r^{\prime}$) defined below.
@@
To evaluate the performance of each type of calculation, we explore the relative reduction in the variational energy with respect to the energy obtained by the calculation of the fixed structure of MPN at $\chi=16$,
$\delta r(\nu,\chi)$ $=$ $\frac{E_{0}(\nu,\chi)-E_{\rm MPN}(\nu,16)}{|E_{\rm MPN}(\nu,16)|},$ (11)
$\delta r^{\prime}(\nu,\chi)$ $=$ $\frac{E^{\prime}_{0}(\nu,\chi)-E^{\prime}_{\rm MPN}(\nu,16)}{|E^{\prime}_{\rm MPN}(\nu,16)|},$ (12)
@@
where $E^{(^{\prime})}_{0}(\nu,\chi)$ is the variational energy obtained at the bond dimension $\chi$ for the random sample $\nu=1,...,\mathcal{N}_{\rm s}$ and $E^{(^{\prime})}_{\rm MPN}(\nu,16)$ is the variational energy of MPN at $\chi=16$. Then, we examine the random averages of $\delta r(\nu,\chi)$ and $\delta r^{\prime}(\nu,\chi)$, respectively denoted as $\delta r(\chi)]$ and $\delta r^{\prime}(\chi)]$. ($\cdots$] denotes the average over $\mathcal{N}_{\rm s}$ random samples.) We note that $\delta r^{(^{\prime})}(\nu,\chi)$ can quantitatively evaluate the degree of reduction in the variational energy $E^{(^{\prime})}(\nu,\chi)$ in each random sample, and therefore, its random average can be a useful indicator of the accuracy improvement of each type of calculation. In the following, we present only the results of $\delta r(\chi)]$ since we have found that $\delta r(\chi)]$ and $\delta r^{\prime}(\chi)]$ lead to essentially the same conclusions.
@@
Figure 4 shows the random average $\delta r(\chi)]$ as a function of $\chi$. Also, the values of $\delta r(\chi=40)]$ for all types are plotted in Fig. 5. The magnitude of the error bar for the random average $\delta r(\chi)]$ shown in the figure is obtained as
$\sigma_{r}=\sqrt{\frac{1}{\mathcal{N}_{\rm s}(\mathcal{N}_{\rm s}-1)}\sum_{\nu=1}^{\mathcal{N}_{\rm s}}\left\{\delta r(\nu,\chi)-[\delta r(\chi)]\right\}^{2}}.$
Figure 4: Random averages of the relative reduction in the variational energy, $\delta r(\chi)]$, for the calculations where the initial TTN is (a) MPN, (b) PBT, and (c) tSDRG-TTN. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. In (a), the data of the calculation without structural optimization for MPN are shown in the inset.
@@
Figure 5: Random averages of the relative reduction in the variational energy, $\delta r(\chi)]$, at $\chi=40$. Circles, squares, and triangles represent the results for the calculations where the initial TTN is MPN, PBT, and tSDRG-TTN, respectively. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the calculations with structural optimization, the results for $\chi_{\rm opt}=16$ ($32$) are shown by solid (open) symbols.
@@
In the results without structural optimization, MPN has the poorest accuracy in variational energies, whereas PBT and tSDRG-TTN yield a relatively good accuracy. As for the results of non-stochastic structural optimization with the bond EE [Eq. (3)] and the truncated rDM weight [Eq. (4)], a significant improvement is observed when the initial TTN is MPN, reflecting the poor accuracy of the MPN calculation without structural optimization. When the initial TTN is PBT or tSDRG-TTN, there is an improvement in accuracy although it is small. Besides, when the initial TTN is MPN and PBT, the accuracy in the variational energies is worse than that of tSDRG-TTN without structural optimization. These results suggest that the results of non-stochastic structural optimization are under the influence of the initial TTN. (Incidentally, this observation indicates the effectiveness of tSDRG for the random XY model.) We note that there is no notable difference between the results of the structural optimizations with the bond EE and the truncated rDM weight; their performance is comparable, at least for the random XY model and $\chi_{\rm opt}$ treated. For the stochastic structural optimization, the variational energy is further reduced, yielding the best accuracy among the results obtained here. Moreover, it is important to note that the results of the stochastic optimization achieve almost the same level of accuracy, independent of the initial TTN. These results suggest that the stochastic structural optimization works well for the random XY model and succeeds in escaping the influence of the initial TTN.
@@
Concerning the effect of $\chi_{\rm opt}$ at which the structural optimization is performed, the results of $\chi_{\rm opt}=32$ yield slightly lower variational energy than those of $\chi_{\rm opt}=16$ for the stochastic optimization. However, the difference is within the statistical error and is basically smaller than the difference from the results of non-stochastic optimizations. This observation indicates that the value of $\chi_{\rm opt}$ has only a minor effect compared to the initial TTN and the updating scheme, at least within the range of $\chi$ we treated here.
@@
Next, we discuss the EEs in the TTN. Since the EE on the physical bonds, the bonds directly connected to the bare spins, takes the same value regardless of the TTN structure, we consider only the EEs on the auxiliary bonds in the following analysis. In the analysis, we use the data of EEs obtained in the sweep which yields the lowest variational energy $E_{0}$ at each $\chi$. For the stochastic structural optimization, we use the EEs obtained by the run that yields the lower variational energy $E_{0}$ at $\chi=40$ between the two runs with different annealing parameters.
@@
Figure 6: (a) $\left[\mathcal{S}_{\rm av}\right]$ and (b) $\left[\mathcal{S}_{\rm max}\right]$, that are respectively the random averages of the average values $\mathcal{S}_{\rm av}(\nu)$ and maximum values $\mathcal{S}_{\rm max}(\nu)$ of EEs on auxiliary bonds in each random sample. Circles, squares, and triangles represent the results at $\chi=40$ for the calculations where the initial TTN is MPN, PBT, and tSDRG-TTN, respectively. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the calculations with structural optimization, the results for $\chi_{\rm opt}=16$ ($32$) are shown by solid (open) symbols. Grey crosses represent the results at $\chi=32$ ($\times$) and $\chi=24$ (++); the latter is not present for the structural optimizations with $\chi_{\rm opt}=32$. The results for the calculation without structural optimization for MPN are plotted in a different scale. Error bars are plotted only for the results at $\chi=40$, although they are smaller than the symbols in most cases.
@@
To see how the EEs in the TTN are suppressed by the structural optimization, we explore the average and maximum values of EE. Figure 6 shows the random averages of $\mathcal{S}_{\rm av}(\nu)$ and $\mathcal{S}_{\rm max}(\nu)$, $\left[\mathcal{S}_{\rm av}\right]$ and $\left[\mathcal{S}_{\rm max}\right]$, where $\mathcal{S}_{\rm av}(\nu)$ and $\mathcal{S}_{\rm max}(\nu)$ are respectively the average and maximum values of EE on the auxiliary bonds in each random sample. Error bars are obtained by a formula similar to Eq. (13). We note that since $\left[\mathcal{S}_{\rm av}\right]$ and $\left[\mathcal{S}_{\rm max}\right]$ are increasing functions of $\chi$, one should analyze them at a sufficiently large $\chi$ to compare the TTNs obtained by different types of structural optimization.
@@
For the average EE, $\left[\mathcal{S}_{\rm av}\right]$, the result for MPN without structural optimization is remarkably large and has a pronounced $\chi$ dependence, while the $\chi$ convergence is almost achieved at $\chi=40$ for the other data. This observation indicates that in MPN, there is a significant fraction of auxiliary bonds whose EEs are too large to be borne even by a bond with $\chi=40$. The non-stochastic structural optimizations with the bond EE and the truncated rDM weight reduce $\left[\mathcal{S}_{\rm av}\right]$ to some extent, but the reductions are modest and their dependence on the initial TTN is visible, indicating that the results are still under the influence of the initial TTNs. In contrast, $\left[\mathcal{S}_{\rm av}\right]$ obtained by the stochastic structural optimization basically achieves small values independently of the initial TTNs. This result demonstrates the effectiveness of the stochastic optimization for the random XY model. As for the dependence of the results on $\chi_{\rm opt}$, the difference between $\left[\mathcal{S}_{\rm av}\right]$ for $\chi_{\rm opt}=16$ and $\chi_{\rm opt}=32$ is basically small compared to the difference due to the initial TTN or the updating scheme, indicating that the effect of $\chi_{\rm opt}$ is secondary.
@@
The data of $\left[\mathcal{S}_{\rm max}\right]$ suffer from relatively large $\chi$ dependence; however, they lead us to the following conclusions that are basically consistent with the ones we read from the data of $\mathcal{S}_{\rm av}]$. That is, the results of the calculation without structural optimization exhibit specific behaviors reflecting the characteristics of each TTN. [We note that $\mathcal{S}_{\rm max}]$ for the fixed structure of MPN at $\chi=40$ is smaller than those for other calculations, but its large $\chi$ dependence implies that it may become greater than the others at $\chi\to\infty$.] The results of the non-stochastic schemes exhibit a sizable dependence on the initial TTN. The results of stochastic structural optimization yield the smallest $\left[\mathcal{S}_{\rm max}\right]$ [except the one for the fixed MPN], with small dependence on the initial TTN. In addition, compared to the differences by the initial TTN and structural optimization scheme, the differences by $\chi_{\rm opt}$ are minor.